# Collaborative Perception Datasets for Autonomous Driving: A Review

### v1

@@ VI Perception Tasks and Metrics
With the development of CP, various datasets have been introduced to support key perception tasks [109][110] . This chapter categorizes them by task type, including object detection, semantic segmentation, object tracking, and trajectory prediction. Object detection focuses on 3D localization of targets, while semantic segmentation enables pixel-level scene understanding. Object tracking and trajectory prediction involve continuous monitoring and future motion inference of dynamic agents. For each task, we present representative datasets and their corresponding evaluation metrics to guide researchers in selecting appropriate datasets and assessment criteria.
@@ VI-A1 Datasets for Object Detection
Multiple datasets support V2X object detection across different collaboration modes and sensing modalities. DAIR-V2X-C [39] provides a real-world V2I benchmark for multi-modal detection. V2X-Sim 2.0 [41], offers 3D detection data for cooperative agents. HoloVIC [85] constructs holographic intersections for multi-view 3D sensing. V2X-Real [54] supports both V2I and V2V detection at scale, while V2V4Real [53] focuses on diverse road types for collaborative 3D detection. OPV2V [40] and OPV2V+ [80] simulate multi-agent cooperation, with the latter offering improved detection benchmarks. For roadside perception, RCooper [60] enhances infrastructure-based 3D detection under occlusions using multiple sensor layouts. Rope3D [75], while not a V2X dataset, provides monocular 3D detection benchmarks that benefit cooperative vision research. Overall, these datasets span real and simulated settings, support V2I and V2V collaboration, and offer diverse modalities, forming a solid foundation for V2X object detection studies.
@@ VI-A2 Metrics for Object Detection
Most CP datasets adopt AP or mAP as core metrics. RCooper [60] reports AP at multiple IoU thresholds. OPV2V [40] and OPV2V+ [80] evaluate vehicle detection at IoU@0.5 and IoU@0.7, with performance also measured within communication ranges. V2X-Real [54] employs both AP and mAP, using lower IoU thresholds (0.3/0.5) to account for varying object sizes. V2X-Sim 2.0 [41] focuses on BEV-based AP at multiple thresholds. Some datasets integrate communication cost into evaluation. V2V4Real [53] reports AP@0.5 and AP@0.7, and introduces Average MegaByte (AM) to quantify transmission overhead under synchronous and asynchronous conditions. DAIR-V2X [39] incorporates communication efficiency and limits evaluation to objects within the ego vehicle’s ROI. Rope3D [75] enhances AP-based evaluation with additional metrics such as Average Ground Center Similarity (ACS), Average Orientation Similarity (AOS), Average Area Similarity (AAS), and Average Four Ground Points Distance (AGD) to assess localization, orientation, and shape accuracy.
In addition to metric design, a range of benchmark methods have been evaluated across typical CP datasets. Table V summarizes the performance of representative algorithms on DAIR-V2X, V2V4Real, OPV2V, and V2XSet, offering insights into algorithm advancements and dataset challenges.
In summary, AP and IoU remain dominant evaluation standards in CP datasets, while extended metrics—such as BEV AP, communication cost, and asynchronous performance—offer deeper insights into real-world constraints. However, AP treats all missed detections equally, regardless of safety relevance. Future research should explore safety-aware evaluation frameworks tailored to CP tasks.
@@ VI-B1 Datasets for Semantic Segmentation
Several datasets support semantic segmentation in CP tasks. V2X-Sim 2.0 [41] provides multi-agent LiDAR segmentation with dynamic obstacle, lane, and BEV semantic annotations. OPV2V [40] supports BEV semantic segmentation using LiDAR-camera fusion, while OPV2V+ [80] enhances cross-agent temporal alignment. OPV2V-N [97] focuses on improving image segmentation robustness. COMAP [79] offers BEV semantic segmentation in simulated urban scenarios. DeepAccident [51] provides accident scene segmentation with fine-grained semantic labels. DOLPHINS [88] supports V2X BEV segmentation under varied weather. Adver-City [50] collects segmentation data under extreme weather using dedicated semantic cameras.
Despite these advances, semantic segmentation in CP still faces several challenges, including inconsistent semantic representations across agents, and reduced robustness in adverse weather or sparse data conditions. Moreover, current datasets are limited in diversity and scale, especially for real-world multi-agent segmentation benchmarks. Future trends will emphasize the construction of large-scale real-world V2X segmentation datasets, cross-agent temporal consistency, and robustness under extreme conditions to support reliable and generalizable perception in complex collaborative scenarios.
@@ VI-B2 Metrics for Semantic Segmentation
Semantic segmentation in CP tasks is commonly evaluated using Pixel Accuracy (PA), mean Pixel Accuracy (mPA), Intersection over Union (IoU), and mean IoU (mIoU) [120, 121]. PA measures the ratio of correctly classified pixels, mPA averages PA across classes to address imbalance, and mIoU reflects overall segmentation accuracy by averaging per-class IoU.
In CP scenarios, evaluation extends beyond single-agent performance to assess cross-agent consistency and modality alignment. V2X-Sim 2.0 [41] evaluates BEV semantic segmentation across categories such as lanes, vehicles, and buildings using mIoU. OPV2V-N [97] focuses on map-level segmentation (drivable area, lanes, vehicles), using IoU as the core metric. Effective CP segmentation evaluation requires traditional metrics and additional measures such as temporal stability, cross-modal IoU, and multi-agent coherence.
@@ VI-C1 Datasets for Tracking and Prediction
Representative datasets include V2X-Seq [42], V2XPnP-Seq [91], and DAIR-V2X-Traj [56]. V2X-Seq offers long-sequence data for both V2V and V2I scenarios, while V2XPnP-Seq enhances temporal consistency and multi-view target association. DAIR-V2X-Traj focuses on V2I trajectory prediction with high-precision annotations. DeepAccident [51] addresses accident prediction by providing trajectory data in critical traffic scenarios.In urban-scale environments, CityFlow [44] supports cross-camera object tracking for traffic monitoring, and INTERACTION [45] provides detailed vehicle trajectories for interactive behavior modeling. RCooper [60] employs roadside sensors to support tracking in dense traffic. OPV2V [40] and V2V4Real [53] benchmark V2V tracking and trajectory prediction in simulated and real-world settings, respectively. V2X-Sim 2.0 [41] supports multi-agent tracking and prediction using BEV-based annotations.
Overall, CP datasets for tracking and prediction span V2V, V2I, and roadside sensing paradigms, providing diverse benchmarks for multi-agent temporal perception and forecasting in AD systems.
@@ VI-C2 Metrics for Tracking and Prediction
Object tracking and trajectory prediction are central to CP, requiring precise evaluation of detection accuracy, tracking stability, and motion forecasting. Standard metrics have been developed to meet these requirements, with extensions introduced to address the complexities of multi-agent V2X scenarios.
For tracking tasks, the CLEAR MOT framework is widely adopted, incorporating Multiple Object Tracking Accuracy (MOTA), Precision (MOTP), and Identity F1 Score (IDF1). MOTA considers False Positives (FP), False Negatives (FN), and Identity Switches (IDSW), while MOTP measures average IoU between predicted and ground-truth trajectories. IDF1 evaluates consistency in identity association. Several datasets extend this framework. RCooper [60] and V2V4Real [53] incorporate Average MOTA (AMOTA), Average MOTP (AMOTP), and scaled AMOTA (sAMOTA) to assess tracking under varied object sizes and detection thresholds. DAIR-V2X-Traj [56] and V2X-Seq [42] also report MOTA, MOTP, and ID Switches for comprehensive analysis. To better reflect multi-agent collaboration, V2X-Sim 2.0 [41] introduces Higher Order Tracking Accuracy (HOTA), which jointly evaluates detection, association, and temporal stability. It also uses Byte Per Second (BPS) to quantify transmission cost, enabling analysis of the trade-off between bandwidth and accuracy.
Trajectory prediction tasks primarily use Average Displacement Error (ADE) and Final Displacement Error (FDE). ADE measures the mean offset over the entire predicted trajectory, while FDE focuses on the final point prediction accuracy. Datasets like INTERACTION [45] and V2X-Seq [42] adopt these metrics, with V2X-Seq further introducing Missing Rate (MR) to capture trajectory coverage completeness. In accident prediction, DeepAccident [51] proposes Volumetric Prediction Quality (VPQ) to jointly evaluate object detection and trajectory forecasting. It also introduces Accident Prediction Accuracy (APA), based on True/False Positives and Negatives, offering a safety-aware evaluation scheme for high-risk scenarios.
Overall, MOTA, MOTP, ADE, and FDE remain core metrics for tracking and prediction, while newer metrics like HOTA, BPS, VPQ, and APA enable more comprehensive evaluations tailored to V2X CP. These frameworks offer critical tools for benchmarking CP performance and guiding future system optimization.
@@ VII-A Availability and Access Methods of Datasets
Dataset accessibility plays a critical role in promoting research progress in CP [123] . Publicly available datasets reduce annotation costs and enable reproducibility, fostering innovation and collaboration. To assess the openness of current datasets, we categorize existing CP datasets based on their access methods.
First, a large portion of datasets can be directly downloaded from official websites, offering high accessibility without requiring registration or application. These include Ko-PER [43], CityFlow [44], LUMPI [76], TUMTraf-I [77], RoScenes [78], OPV2V [40], OPV2V+ [80], V2V4Real [53], LUCOOP [81], TUMTraf V2X [87], V2X-Sim 2.0 [41], DeepAccident [51], V2X-Real [54], Multi-V2X [89], Adver-City [50], MARS [82], Mixed Signals [92], and SCOPE [93]. In addition, some datasets are hosted on GitHub, often accompanied by source code, benchmark tools, or baseline implementations. Representative examples include V2V-Sim [47], OPV2V-H [83], CoPerception-UAV [35], CoopInf [48], V2XSet [34], and V2X-R [58], which facilitate rapid experimentation and method replication. Meanwhile, several datasets are distributed via cloud platforms such as Baidu Netdisk, Google Drive, or Alibaba Cloud. Access typically requires a request or registration process. These include H-V2X [52], DAIR-V2X-C [39], V2X-Seq [42], OTVIC [86], V2XReid [55], V2X-Radar [57], WHALES [90], and V2XPnP-Seq [91]. Furthermore, a number of datasets require formal application procedures, often involving submission of research proposals or institutional endorsements. Examples include BAAI-VANJEE [108], Rope3D [75], A9-Dataset [73], HoloVIC [85], V2V-QA [84], CODD [49], DAIR-V2X-Traj [56], INTERACTION [45], IPS300+ [74], DOLPHINS [88], InScope [59], and RCooper [60]. Finally, a few datasets do not publicly disclose their access procedures, possibly due to ongoing internal usage, limited availability, or institutional restrictions. These include WIBAM [105], T & J [46], IRV2V [98], OPV2V-N [97], CARTI [124], and V2X-DSI [125].
In summary, while many CP datasets are openly accessible via official websites, GitHub, or cloud storage, a significant portion remains gated behind application-based processes or lacks clear documentation. Improving data openness, unifying access procedures, and providing transparent distribution mechanisms will be essential to advancing CP research.
@@ VII-B Comprehensive Quality Analysis of Datasets
Overall, high-quality datasets exhibit complementary characteristics: real-world datasets like DAIR-V2X [39] and V2X-Real [54] emphasize fidelity and generalization, while simulation-based datasets like OPV2V+ [80] and Multi-V2X [89] excel in scale, flexibility, and controlled benchmarking. The increasing inclusion of diverse modalities, rich annotations, and comprehensive task coverage highlights a clear trend toward more holistic, multi-agent CP datasets. Future research would benefit from standardized benchmarking protocols across these high-quality datasets to promote reproducibility and cross-domain evaluation.
@@ VII-C Summary of High-Influence Datasets
High-impact datasets are typically defined by high citation rates and widespread use in benchmark evaluations. For infrastructure-based perception, CityFlow [44], INTERACTION [45], and Rope3D [75] are widely used in multi-camera tracking and urban scene understanding. Representative methods such as BEVHeight [126], BEVHeight++ [127], and ViT-CoMer [128] have been developed using these datasets. In V2V perception, OPV2V [40] is one of the most cited simulation datasets, supporting algorithms like Where2comm [35], CoBEVT [114], and OPV2V-H [83]. In the V2I domain, DAIR-V2X [39] and its extensions (e.g., DAIR-V2X-Seq [42], DAIR-V2X-Traj [56]) provide large-scale real-world data with high-quality multimodal annotations, supporting tasks such as detection, tracking, and motion forecasting. These datasets underpin studies including CoCa3D [80], HM-ViT [129], BEV-V2X [130], and How2comm [131]. In the V2X category, V2X-Sim [41] supports BEV detection and semantic segmentation, and is frequently adopted by methods such as VoxFormer [132] and V2X-ViT [34].
@@ VIII-A4 Datasets for Extreme Scenarios
Extreme scenarios are critical for evaluating the safety of AD systems. However, existing datasets contain few extreme scenario samples, limiting model generalization in these key situations. For example, the DeepAccident dataset attempts to provide traffic accident data, but its coverage of accident types remains incomplete. Future research should explore generating extreme scenario data through simulation, combined with real-world data augmentation, to enhance model responsiveness in hazardous situations. Moreover, anomaly detection and Out-Of-Distribution (OOD) analysis can help identify extreme scenarios, improving the safety and stability of CP systems.
@@ VIII-D Domain Adaptation and Generalization
Domain adaptation remains a key challenge in improving dataset generalization. Due to differences in data collection environments, sensor configurations, and traffic scenarios, models trained on one dataset may not perform well on another. For instance, CP models trained in simulation environments often struggle when applied to real-world scenarios. Future research should explore unsupervised domain adaptation and self-supervised learning methods to enhance cross-domain generalization. Additionally, using Generative Adversarial Networks (GANs) to synthesize cross-domain data and leveraging large-scale pre-trained models for domain transfer learning could provide effective solutions.
@@ IX Conclusion
This review presents a comprehensive review of CP datasets for AD, classifying them by collaboration paradigms, sensor modalities, data collection methods, environmental scenarios, and supported tasks. We analyze both real-world and simulation-based datasets across V2V, V2I, V2X, and I2I settings, evaluating their annotation quality, accessibility, scenario coverage, and benchmarking readiness. While existing datasets have significantly advanced CP research, limitations remain in data scale, coverage of rare and adverse conditions, agent diversity, and geographic generalization. To bridge these gaps, future datasets should incorporate long-tail scenarios such as traffic anomalies and extreme weather, include data from diverse regions, support heterogeneous agent types, and reflect realistic communication constraints. Establishing standardized annotation formats and unified evaluation protocols is essential for enabling fair comparisons and fostering cross-dataset generalization. Additionally, techniques such as automated labeling, high-fidelity simulation, and generative modeling—especially with the aid of large language models—offer promising directions for building scalable and semantically rich datasets. We hope this review provides valuable guidance for the design, construction, and benchmarking of future datasets, ultimately advancing CP systems in real-world AD applications.

### v2

@@ VI Perception Tasks and Metrics
With the development of CP, various datasets have been introduced to support key perception tasks [116][117] . This chapter categorizes them by task type, including object detection, semantic segmentation, object tracking, and trajectory prediction. Object detection focuses on 3D localization of objects, while semantic segmentation enables pixel-level scene understanding. Object tracking and trajectory prediction involve continuous monitoring and future motion inference of dynamic agents. For each task, we present representative datasets and their corresponding evaluation metrics to guide researchers in selecting appropriate datasets and assessment criteria.
@@ VI-A1 Datasets for Object Detection
Numerous datasets have been developed to support V2X object detection across varying collaboration paradigms and perception modalities. Real-world datasets such as DAIR-V2X-C [35], V2X-Real [53], and V2V4Real [52] provide large-scale, multi-modal annotations for V2I and V2V scenarios. Simulated datasets like OPV2V [39], OPV2V+ [81], and V2X-Sim 2.0 [40] offer controlled environments for benchmarking. HoloVIC [86] and RCooper [60] emphasize occlusion-aware perception via infrastructure-side perception. Although Rope3D [76] is not V2X-specific, it contributes monocular 3D benchmarks relevant to cooperative perception. Collectively, these datasets form a diverse foundation for V2X detection research.
As shown in Fig. 11, recent methods demonstrate improved performance across benchmarks. On V2V4Real, ERMVP [118] produces more complete and better-aligned bounding boxes than V2VNet [46] and CoBEVT [119], especially in dense scenes. On DAIR-V2X, V2X-ViTv2 [120] detects more dynamic vehicles with lower spatial error than V2VNet and Where2comm [34], aided by transformer-based fusion. On OPV2V, MKD-Cooper [121] surpasses DiscoNet [122] and V2X-ViT [33] in producing orientation-aware results with fewer false positives on complex road geometries. These comparisons highlight the impact of advanced fusion strategies in both real and simulated domains.
Figure 11: Qualitative detection results across three datasets: (a–c) V2V4Real with V2VNet, CoBEVT, and ERMVP [118]; (d–f) DAIR-V2X with V2VNet, Where2comm, and V2X-ViTv2 [120]; (g–i) OPV2V with OPV2V, DiscoNet, and MKD-Cooper [121]. Green boxes show ground truth; red/orange boxes indicate predictions. ERMVP and MKD-Cooper yield more accurate and robust results in complex scenes. Images reproduced from the original publications with proper citation.
@@ VI-A2 Metrics for Object Detection
Most datasets use AP or mAP as core metrics. RCooper [60] reports AP under multiple IoUs; OPV2V [39] and OPV2V+ [81] assess detection at 0.5 and 0.7 IoU, including within communication range. V2X-Real [53] adopts lower thresholds (0.3/0.5) for scale variance. V2X-Sim 2.0 [40] emphasizes BEV-based AP. V2V4Real [52] introduces Average MegaByte (AM) to quantify communication cost under sync/async settings, while DAIR-V2X [35] evaluates within ego-ROI under bandwidth constraints. Rope3D [76] supplements AP with metrics like ACS, AOS, AAS, and AGD for comprehensive localization and shape assessment.
Beyond metrics, representative methods are benchmarked on datasets such as DAIR-V2X, V2V4Real, OPV2V, and V2XSet, as summarized in Table V, revealing advances and ongoing challenges. In conclusion, AP and IoU remain central for evaluating CP detection, while extended metrics—e.g., BEV AP, communication overhead, and asynchronous accuracy—offer deeper realism. Future work should incorporate safety-aware evaluation to prioritize critical perception failures.
@@ VI-B1 Datasets for Semantic Segmentation
Several datasets support semantic segmentation for CP. V2X-Sim 2.0 [40] provides multi-agent LiDAR segmentation with BEV and dynamic obstacle annotations. OPV2V [39] enables BEV segmentation via LiDAR-camera fusion, while OPV2V+ [81] and OPV2V-N [99] improve temporal alignment and image-based robustness, respectively. COMAP [80] offers simulated urban BEV annotations. DeepAccident [50] provides fine-grained accident scene labels. DOLPHINS [89] and Adver-City [49] support segmentation under adverse conditions using diverse sensor setups.
Challenges remain, including inconsistent semantic labeling across agents and limited robustness under sparse or extreme conditions. Real-world datasets are also limited in scale and diversity. Future efforts should focus on large-scale, real-world, multi-agent segmentation benchmarks with strong cross-agent consistency and resilience to weather and domain shifts.
@@ VI-B2 Metrics for Semantic Segmentation
Semantic segmentation in CP is typically evaluated using Pixel Accuracy (PA), mean PA (mPA), Intersection over Union (IoU), and mean IoU (mIoU) [130][131]. PA measures the proportion of correctly classified pixels, while mPA and mIoU address class imbalance and overall accuracy. In CP settings, evaluation extends to inter-agent consistency and modality fusion. V2X-Sim 2.0 [40] reports mIoU for semantic categories such as lanes and vehicles. OPV2V-N [99] focuses on map-level segmentation using IoU. Advanced evaluation should also consider temporal stability, cross-modal consistency, and multi-agent alignment.
@@ VI-C1 Datasets for Tracking and Prediction
Representative datasets for tracking and prediction span various CP paradigms. V2X-Seq [41] and V2XPnP-Seq [92] support long-sequence V2V/V2I perception with improved temporal consistency and multi-view association. DAIR-V2X-Traj [55] provides accurate V2I trajectory annotations, while DeepAccident [50] focuses on forecasting in safety-critical scenarios. Urban-scale datasets like CityFlow [43] and INTERACTION [44] enable cross-camera tracking and behavior modeling. RCooper [60] captures dense roadside tracking scenes, and OPV2V [39] and V2V4Real [52] benchmark V2V tracking in simulation and real-world contexts. V2X-Sim 2.0 [40] supports BEV-based multi-agent prediction. Collectively, these datasets provide diverse benchmarks for evaluating temporal perception across V2V, V2I, and infrastructure-based CP.
@@ VI-C2 Metrics for Tracking and Prediction
Tracking and prediction in CP rely on metrics that evaluate detection accuracy, identity consistency, and MF. Standard CLEAR MOT metrics—MOTA, MOTP, and IDF1—remain widely used. Datasets like RCooper [60] and V2V4Real [52] extend these with AMOTA and sAMOTA for scale-aware evaluation. DAIR-V2X-Traj [55] and V2X-Seq [41] report similar metrics, while V2X-Sim 2.0 [40] introduces HOTA and BPS to account for association quality and communication cost. Trajectory prediction is typically assessed using ADE and FDE, as adopted by INTERACTION [44] and V2X-Seq [41], with the latter also reporting Missing Rate (MR). DeepAccident [50] introduces VPQ and APA to evaluate detection-prediction synergy and accident classification. While MOTA and ADE remain core metrics, scenario-aware metrics like HOTA and VPQ offer deeper performance insights.
@@ VII-A Availability and Access Methods of Datasets
Dataset accessibility plays a crucial role in advancing CP research [133]. Public availability reduces annotation costs, enhances reproducibility, and promotes innovation across the community. To evaluate current openness, we categorize CP datasets by their access methods.
Many datasets, including Ko-PER [42], CityFlow [43], OPV2V [39], V2V4Real [52], and DeepAccident [50], support direct download from official websites without registration. Others, such as V2V-Sim [46], V2XSet [33], and CoopInf [47], are hosted on GitHub alongside code and benchmarks, facilitating reproducible research. Datasets like DAIR-V2X-C [35], OTVIC [87], and WHALES [91] require requests through cloud storage services. Some datasets, including Rope3D [76], INTERACTION [44], and RCooper [60], involve formal application processes and institutional approval. A few others, such as IRV2V [100] and OPV2V-N [99], lack clear access instructions. Detailed access types and download links are summarized in Table II.
In summary, while many CP datasets are readily available, a notable portion remains behind restricted access or insufficient documentation. Enhancing openness, standardizing access procedures, and improving transparency in dataset distribution will be key to accelerating progress in CP research.
@@ VII-B Comprehensive Quality Analysis of Datasets
High-quality CP datasets exhibit complementary strengths: real-world datasets such as DAIR-V2X [35] and V2X-Real [53] prioritize realism and generalization, while simulation datasets like OPV2V+ [81] and Multi-V2X [90] offer scalability and control. The trend toward multimodal data, rich annotations, and task diversity suggests a shift to more holistic, multi-agent CP benchmarks. Standardized evaluation protocols are needed to enhance reproducibility and cross-domain comparability.
@@ VII-C Summary of High-Influence Datasets
High-impact datasets are characterized by high citation rates and widespread benchmark adoption. In infrastructure-based perception, CityFlow [43], INTERACTION [44], and Rope3D [76] are widely used for multi-camera tracking and urban understanding, supporting methods like BEVHeight [134], BEVHeight++ [135], and ViT-CoMer [136]. In V2V, OPV2V [39] enables numerous algorithms including Where2comm [34], CoBEVT [119], and OPV2V-H [84]. DAIR-V2X [35] and its variants [41, 55] support V2I tasks like detection and tracking, used in CoCa3D [81], HM-ViT [137], and BEV-V2X [138]. For V2X, V2X-Sim [40] facilitates BEV detection and segmentation, widely adopted by VoxFormer [139] and V2X-ViT [33].
@@ VIII-A4 Datasets for Extreme Scenarios
Extreme scenarios are critical for evaluating the safety of AD systems. However, existing datasets contain few extreme scenario samples, limiting model generalization in these key situations. For example, the DeepAccident dataset attempts to provide traffic accident data, but its coverage of accident types remains incomplete. Future research should explore generating extreme scenario data through simulation, combined with real-world data augmentation, to enhance model responsiveness in hazardous situations. Moreover, anomaly detection and Out-Of-Distribution analysis can help identify extreme scenarios, improving the safety and stability of CP systems.
@@ VIII-D Domain Adaptation and Generalization
Domain adaptation is a core challenge for improving model generalization across datasets. Differences in sensors, environments, and traffic patterns often degrade performance when models are transferred across domains. In particular, CP models trained on simulation data struggle in real-world deployment due to noise, occlusion, and latency mismatches. Future research should explore unsupervised and self-supervised methods to bridge this Sim2Real gap. Promising directions include domain-aligned feature learning, simulation-aware augmentation, and leveraging large-scale pre-trained models. Additionally, Generative Adversarial Networks can help synthesize realistic cross-domain samples to support robust transfer.
@@ IX Conclusion
This review provides a comprehensive overview of CP datasets for AD, categorized by collaboration paradigms, sensor modalities, data sources, environments, and supported tasks. We examine real and simulated datasets across V2V, V2I, V2X, and I2I settings, assessing their annotation quality, accessibility, scenario diversity, and benchmarking readiness. Despite notable progress, challenges remain in data scale, rare condition coverage, agent diversity, and regional generalization. Future datasets should address long-tail scenarios, broaden geographic diversity, support heterogeneous agents, and capture realistic communication constraints. Standardized annotations and unified evaluation protocols are critical for fair comparison and cross-dataset generalization. Advances such as automated labeling, high-fidelity simulation, and LLM-assisted generative modeling offer promising pathways for scalable, semantically rich datasets. We hope this review informs future dataset design and accelerates CP research toward robust real-world AD applications.